OpenStudy (anonymous):

greg noticed the four fractions below in the mathematics book 2/5, 3/8, 4/8, 1/6 what are they least to greatest please help me

OpenStudy (anonymous):

You have to get a common denominator in order to compare these numbers. So with 2/5, 3/8, 4/8, and 1/6, you have to find a number 5, 8, and 6 can be multiplied to get. Honestly, I cant even think of the lowest common denominator in this case, so I just multiplied 8x6x5 to get 240 Now multiply each fraction by whatever number it takes to get 240, remembering that whatever you do to the top number, you have to do to the bottom to keep it equivalent. For 2/5, multiply both top and bottom by 48 = 96/240 For 3/8, multiply both top and bottom by 30 = 90/240 For 4/8, multiply both top and bottom by 30 = 120/240 For 1/6 multiply both top and bottom by 40 = 40/240 Now just rank by the numerator, top numbers so 4/8 > 2/5> 3/8> 1/6

OpenStudy (anonymous):

thank you im really horrible in math but somehow im in gifted classes so im in pre algebra this year and i really needed help thank you

OpenStudy (anonymous):

No problem, notice though that I listed them greatest to least.